OpenStudy (anonymous):

use the lim f(x)-f(a)/x-a to find the derivative x->a f(x)=-3x^2-5x+1 and point (1,-7)

OpenStudy (anonymous):

I know how to set it up can't factor it

OpenStudy (anonymous):

so lim(x->a) (3x²-5x-3a²+5a)/(x-a)?

OpenStudy (anonymous):

it has to factor because 1 will give you a zero in the numerator. so it must factor as \[(x-1)(\text{something})\]

OpenStudy (anonymous):

They want to use the limit definition to find the derivative. even though we know the power rule

OpenStudy (anonymous):

ok the idea is this. you know \[f(a)=f(1)=-7\] so \[\frac{f(x)-f(a)}{x-a}=\frac{f(x)-(-7)}{x-1}=\frac{-x^2-5x+1+7}{x-1}=\frac{-x^2-5x+8}{x-1}\]

OpenStudy (anonymous):

except i made a typo so let me fix it

OpenStudy (anonymous):

\[\frac{f(x)-f(a)}{x-a}=\frac{f(x)-(-7)}{x-1}=\frac{-3x^2-5x+1+7}{x-1}=\frac{-3x^2-5x+8}{x-1}\]

OpenStudy (anonymous):

right I got up to that part but stuck here

OpenStudy (anonymous):

you know for a fact that if you replace x by 1 in the numerator you will get zero. that has to be because you subtracted f(1)

OpenStudy (anonymous):

this means it has to factor, and it has to factor as \[(x-1)(\text{whatever})\]

OpenStudy (anonymous):

3x²-5x-3a²+5a factors to 3(x²-a²)-5(x-a)=(x-a)(3(x+a)-5)

OpenStudy (anonymous):

so factor out an x -1 and see what the other term has to be. it has to be \[\frac{(x-1)(-3x-8)}{x-1}\]

OpenStudy (anonymous):

now cancel the factor of x - 1 to get \[-3x-8\] and finally replace x by 1 to get \[-11\]

OpenStudy (anonymous):

oh no, it is right

OpenStudy (anonymous):

could you show me how you factored the top

OpenStudy (anonymous):

sure, but lets make sure one thing is clear. you have the numerator of \[-3x^2-5x+8\] right? and how did you get it? you took \[-3x^2-5x+1\] replaced x by 1 and got -7, then subtracted it off to get \[-3x^2-5x+8\]

OpenStudy (anonymous):

which means for sure if you replace x by 1 you have to get 0

OpenStudy (anonymous):

so you already have one zero of this polynomial, it is 1. which means it MUST factor as \[(x-1)(\text{something})\]

OpenStudy (anonymous):

your only job now is to figure out what the something is. so it is in fact much easier than factoring in general because you already know what one of the factors must be. it must be (x- 1)

OpenStudy (anonymous):

so ask yourself if \[-3x^2-5x+8=(x-1)(\text{something})\] what must the something be? well the first term must be -3x and the second term must be -8 or else it wont work

OpenStudy (anonymous):

so is it safe to assume that the denominator is going to be a factor (in this x-1)

OpenStudy (anonymous):

it is safe to assume that if you replace x by 1 you get zero so it MUST factor that way, yes

OpenStudy (anonymous):

that is what the "factor" theorem says. if you have a zero of a polynomial you can factor it out. so don't go looking to factor as in "gee what can it be" you already know one of the factors, you only have to find the other one

OpenStudy (anonymous):

and certainly it will cancel with the denoninator. of course you already know the derivative is \[f'(x)=-6x-5\] and if you replace x by 1 here you will get \[f'(1)=-11\]

OpenStudy (anonymous):

I get the rest just got stuck on the factoring thanks a lot